# Multimodal Chain of Continuous Thought for Latent-Space Reasoning in Vision-Language Models

### v1

@@ Abstract
Many reasoning techniques for large multimodal models adapt language model approaches, such as Chain-of-Thought (CoT) prompting, which express reasoning as word sequences. While effective for text, these methods are suboptimal for multimodal contexts, struggling to align audio, visual, and textual information dynamically. To explore an alternative paradigm, we propose the Multimodal Chain of Continuous Thought (MCOUT), which enables reasoning directly in a joint latent space rather than in natural language. In MCOUT, the reasoning state is represented as a continuous hidden vector, iteratively refined and aligned with visual and textual embeddings, inspired by human reflective cognition. We develop two variants: MCOUT-Base, which reuses the language model’s last hidden state as the continuous thought for iterative reasoning, and MCOUT-Multi, which integrates multimodal latent attention to strengthen cross-modal alignment between visual and textual features. Experiments on benchmarks including MMMU, ScienceQA, and MMStar show that MCOUT consistently improves multimodal reasoning, yielding up to 8.23% accuracy gains over strong baselines and improving BLEU scores up to 8.27% across multiple-choice and open-ended tasks. These findings highlight latent continuous reasoning as a promising direction for advancing LMMs beyond language-bound CoT, offering a scalable framework for human-like reflective multimodal inference. Code is available at https://github.com/Hanhpt23/OmniMod.
@@ 1 Introduction
Vision-language models (VLMs) have transformed multimodal tasks, such as visual question answering (VQA), image captioning, and reasoning on benchmarks like ScienceQA Lu et al. (2022), MMMU Yue et al. (2024), and IQBench Pham et al. (2025a), by seamlessly integrating visual and textual data. These models leverage visual models and large language models (LLMs) to process heterogeneous inputs, enabling applications from autonomous systems to interactive assistants. However, achieving robust reasoning in VLMs remains a challenge due to limitations in existing techniques, such as attention mechanism within the transformer decoder Vaswani et al. (2017), prompting strategies like CoTs Wei et al. (2022); Yao et al. (2023); Besta et al. (2024), or training methods like reinforcement learning (RL) Ouyang et al. (2022); Pham and Ngo (2025). CoT, originally developed for LLMs, prompts models to generate intermediate reasoning steps in natural language, while other approaches, such as fine-tuning with visual-text alignment, aim to enhance multimodal reasoning. These methods often rely on discrete token sequences or static vision features, leading to computational inefficiencies and difficulties in dynamically aligning visual and textual modalities for coherent reasoning, particularly in tasks requiring fine-grained multimodal understanding.
Inspired by human cognition, where reasoning involves generating intermediate thoughts and iteratively validating them against input data, such as revisiting images or documents to ensure coherence, we propose the Multimodal Chain of Continuous Thought (MCOUT), a novel framework for efficient reasoning in VLMs. MCOUT operates in a unified latent space, dynamically aligning visual and textual representations to mimic reflective human thinking. We introduce two variants: MCOUT-Base, which uses the language model’s last hidden state as a continuous thought for iterative refinement inspired from COCONUT Hao et al. (2024), and MCOUT-Multi, which enhances cross-modal alignment by combining the hidden state with input embeddings via a multimodal latent attention mechanism. By overcoming the limitations of token-based CoT and static vision features, MCOUT reduces computational overhead by directly feeding hidden layers, with/without input embeddings, into the model as continuous thoughts. Tested on diverse benchmarks, MCOUT achieves significant performance gains, positioning it as a pioneering advancement in vision-language reasoning and offering a scalable approach for robust multimodal inference.
@@ 2 Literature Review
The development of reasoning capabilities in VLMs is critical for tasks like VQA and multimodal reasoning. Over the past few years, various techniques have been explored to enhance reasoning in both LLMs and VLMs, including attention mechanism, prompting techniques, training methods, and recent latent reasoning paradigms. CoT prompting, introduced by Wei et al. (2022), has significantly improved LLM performance on arithmetic tasks (e.g., GSM8K) and logical reasoning tasks (e.g., AQUA-RAT) by generating explicit intermediate steps. Building on CoT, its variants have emerged to address reasoning limitations. Self-consistency (Wang et al., 2022) samples multiple CoT outputs and selects the most consistent answer via majority voting, enhancing robustness but increasing computational cost. Tree of Thoughts (ToT) (Yao et al., 2023) structures reasoning as a tree search, exploring multiple paths for complex problem-solving, though its token-based nature remains computationally intensive. Graph of Thoughts (GoT) (Besta et al., 2024) extends ToT by modeling reasoning as a graph, enabling dynamic recombination of thoughts for greater efficiency. In the VLM domain, Multimodal CoT (Zhang et al., 2023) generates interleaved text and image reasoning steps, improving performance on ScienceQA but struggling with cross-modal alignment due to reliance on static vision features and verbose token sequences. These CoT methods, while effective for LLMs, face challenges in VLMs, where aligning heterogeneous modalities and minimizing token overhead are critical.
@@
Beyond prompting, training techniques have been pivotal in enhancing reasoning for both LLMs and VLMs. RL methods, such as those explored by Ouyang et al. (2022), optimize LLMs using human feedback to improve reasoning and alignment, as seen in models like InstructGPT. Group relative policy Optimization (GRPO) (Shao et al., 2024) refines model outputs by incorporating reward signals, enhancing performance on complex tasks. Reasoning functions, such as RARL Pham and Ngo (2025), enable models to learn structured reasoning patterns through optimization, improving logical consistency. RL-based fine-tuning (Shen et al., 2025) has been applied to align visual and textual features, though these methods often rely on static embeddings, limiting dynamic reasoning capabilities. These training techniques complement prompting but still face challenges in efficiently integrating multimodal data for coherent reasoning.
@@
To overcome these limitations, latent reasoning paradigms have shifted reasoning from discrete token sequences to continuous latent spaces. COCONUT Hao et al. (2024) leverages the last hidden state of an LLM as a "continuous thought," enabling parallel exploration of reasoning paths via breadth-first search. This approach reduces token overhead and outperforms CoT on tasks requiring backtracking. Other latent reasoning methods for LLMs include Latent Reasoning Skills (LaRS) Xu et al. (2023), which uses unsupervised learning to create latent representations of rationales, selecting in-context learning examples based on reasoning skills and achieving fourfold faster processing than CoT. Similarly, Wang et al. (2025) proposed a recurrent depth approach that iteratively refines latent representations, scaling test-time computation to enhance performance on complex tasks. These methods demonstrate the efficiency of latent reasoning in LLMs but are primarily designed for text-only contexts, leaving their application to VLMs largely unexplored.
@@
In the VLM domain, latent reasoning is an emerging area with promising developments. Zhang et al. (2023) introduced a multimodal CoT framework that uses diffusion processes to learn a text-image aligned latent space, generating dynamic image features that improve reasoning on ScienceQA and multimodal machine translation. Yang et al. (2025a) developed MMaDA, a diffusion-based VLM that operates in latent spaces for coherent generation and reasoning across text and images, achieving strong performance in tasks like VQA and image captioning. Yang et al. (2025b) proposed the Mirage framework, which augments VLMs with latent visual tokens during decoding, enhancing reasoning efficiency in complex multimodal tasks. Fan and Zhou (2018) introduced stacked latent attention, preserving spatial information in latent spaces to improve reasoning in VQA tasks. Recent efforts, such as multimodal latent language modeling (Sun et al., 2024), employ next-token diffusion for continuous reasoning, while Corvid (Jiang et al., 2025) and Grounded Chain-of-Thought (GCoT) (Wu et al., 2025) address visual hallucination and decision-making accuracy. Despite these advances, most approaches rely on discrete token-based reasoning or static vision features, limiting efficient cross-modal alignment.
@@
Inspired by human cognition and previous work Hao et al. (2024), where reasoning involves generating intermediate thoughts and iteratively validating them against input data, our MCOUT addresses these gaps by introducing a novel latent reasoning framework for VLMs, specifically for image-based tasks. MCOUT employs two variants: MCOUT-Base, which uses the language model’s last hidden state as a continuous thought for iterative refinement, and MCOUT-Multi, which integrates the hidden state with image embeddings via a multimodal latent attention mechanism, enabling dynamic alignment of visual and textual representations. MCOUT mimics human reflective reasoning by iteratively refining thoughts in a continuous latent space, as demonstrated in our implementation, which supports multimodal inputs and has been tested successfully for vision-language reasoning. MCOUT offers a significant advancement, bridging the efficiency of latent reasoning with the complexity of vision-language reasoning, paving the way for robust and scalable VLMs.
@@ 3.1 Model Architecture
The MCOUT framework is built upon a vision-language model, SilVar Pham et al. (2025b), comprising a pre-trained visual encoder $\mathcal{V}$ and a language model $\mathcal{L}$, as illustrated in Figure 1. We use CLIP Radford et al. (2021) as the visual encoder $\mathcal{V}$, which processes input images $\mathbf{x}_{v}\in\mathbb{R}^{H\times W\times C}$ to produce visual embeddings $\mathbf{e}_{v}\in\mathbb{R}^{S_{v}\times D}$, where $S_{v}$ is the sequence length of visual tokens and $D$ is the embedding dimension. For the language model $\mathcal{L}$, we employ Llama 3.2 1B, which processes tokenized text inputs $\mathbf{x}_{t}$ to generate contextual embeddings $\mathbf{e}_{t}\in\mathbb{R}^{S_{t}\times D}$, where $S_{t}$ is the sequence length of text tokens. In this study, we use CLIP and Llama 3.2 1B for all experiments because we want to focus on latent reasoning for small VLMs, although our pipeline is compatible with other LLMs.
@@ 4.1 Datasets and Training
To evaluate the effectiveness of our MCOUT framework, we conducted experiments using four diverse vision-language datasets: VQAv2 Goyal et al. (2017), MMMU Yue et al. (2024), ScienceQA Lu et al. (2022), and MMStar Chen et al. (2024). These datasets assess the model’s reasoning capabilities across multimodal tasks, including VQA, scientific reasoning, and general knowledge understanding, with a focus on image-text integration. The VQAv2 dataset, used for pretraining, contains 443,757 question-answer pairs associated with images from the COCO dataset, emphasizing tasks like object recognition, attribute identification, and spatial reasoning.
@@ 4.2 Results and Benchmarking
Kosmos2 Peng et al. (2023) 1.7 32.70
SilVar Pham et al. (2025b) 7 63.21
LLaVA-7B Liu et al. (2023) 7 41.10
InstructBLIP-7B Dai et al. (2023) 8 54.10
OpenFlamingo Awadalla et al. (2023) 9 44.80
Qwen-VL Bai et al. (2023) 9.6 61.10
MiniGPT-4 Zhu et al. (2023) 13 47.71
LLaMA2-13B Yang et al. (2023) 13 55.78
LLaVA-13B Yang et al. (2023) 13 47.74
PandaGPT-13B Su et al. (2023) 13 63.20
@@
Models Parameters (B) accuracy (%) BLEU
Our experiments
Baseline 1 25.44 25.44
MCOUT-Base ($N_{t}$ = 5) 1 27.53 ($\uparrow$ 8.21%) 27.54 ($\uparrow$ 8.31%)
MCOUT-Multi ($N_{t}$ = 5) 1 27.18 ($\uparrow$ 6.79%) 27.19 ($\uparrow$ 6.82%)
MCOUT-Base ($N_{t}$ = 10) 1 27.52 ($\uparrow$ 8.18%) 27.54 ($\uparrow$ 8.31%)
MCOUT-Multi ($N_{t}$ = 10) 1 27.36 ($\uparrow$ 7.54%) 27.37 ($\uparrow$ 7.58%)
Literature reports
Kosmos 2 Peng et al. (2023) 1.7 23.7
MiniGPT-4-v1-7B Zhu et al. (2023) 7 23.6
LLaVA-v1.5-7B Liu et al. (2023) 7 33.7
MiniGPT-4-v2 Chen et al. (2023) 7 25.0
OpenFlamingo v2 Awadalla et al. (2023) 9 28.8
Qwen-VL Bai et al. (2023) 9.6 29.6
LLaVA-v1.5-13B Liu et al. (2023) 13 37.0
PandaGPT-13B Su et al. (2023) 13 32.9
@@
Models Parameters (B) accuracy (%) BLEU
Our experiments
Baseline 1 25.13 25.14
MCOUT-Base ($N_{t}$ = 10) 1 26.13 ($\uparrow$ 3.98%) 26.14 ($\uparrow$ 3.98%)
MCOUT-Multi ($N_{t}$ = 10) 1 26.07 ($\uparrow$ 3.74%) 26.08 ($\uparrow$ 3.74%)
Literature reports
Kosmos2 Peng et al. (2023) 1.7 24.9
MiniGPT-4-v1-7B Zhu et al. (2023) 7 16.3
MiniGPT-4-v2 Chen et al. (2023) 7 21.3
LLaVA-7B Liu et al. (2023) 7 27.1
OpenFlamingo v2 Awadalla et al. (2023) 9 26.9
Qwen-VL-Chat Bai et al. (2023) 9.6 34.5
PandaGPT-13B Su et al. (2023) 13 25.6
@@ 4.3 Multimodal Latent Reasoning Analysis
MCOUT-Multi, which integrates the last hidden state with multimodal input embeddings, shows last hidden layer means ranging from -0.02212 to -0.00112 across iterations, with standard deviations around 2.24, closely tracking MCOUT-Base’s patterns and indicating minimal disruption from multimodal fusion. However, the mixed embeddings reveal a critical limitation: their mean remains constant at 0.002418 with a negligible standard deviation of 0.001866, and the mixed standard deviation is uniformly 0.198925 across all iterations, suggesting a static, low-variance contribution from the multimodal input. This persistent uniformity, despite normalization, points to a modality collapse, where the attention mechanism fails to extract diverse visual context, aligning MCOUT-Multi’s performance (58.45% accuracy at $N_{t}=5$) closely with MCOUT-Base (58.60%). This observation resonates with the sinking of visual attention in recent studies Kang et al. (2025); Cancedda (2024); Sim et al. (2025), which attributes such collapse to activation imbalances favoring a specific type (e.g. text). Our study shows that low-variance embeddings (mixed_embeds std $\approx$ 0.2 vs. last_hidded std $\approx$ 2.2) limit multimodal benefits. The pre-training norm adjustment likely prevented catastrophic fusion failure, but the static mixed embeddings suggest entropy collapse Zhai et al. (2023), where uniform attention weights diminish multimodal impact.

### v2

@@ Abstract
Many reasoning techniques for large multimodal models adapt language model approaches, such as Chain-of-Thought (CoT) prompting, which express reasoning as word sequences. While effective for text, these methods are suboptimal for multimodal contexts, struggling to align audio, visual, and textual information dynamically. To explore an alternative paradigm, we propose the Multimodal Chain of Continuous Thought (MCOUT), which enables reasoning directly in a joint latent space rather than in natural language. In MCOUT, the reasoning state is represented as a continuous hidden vector, iteratively refined and aligned with visual and textual embeddings, inspired by human reflective cognition. We develop two variants: MCOUT-Base, which reuses the language model’s last hidden state as the continuous thought for iterative reasoning, and MCOUT-Multi, which integrates multimodal latent attention to strengthen cross-modal alignment between visual and textual features. Experiments on benchmarks including MMMU, ScienceQA, and MMStar show that MCOUT consistently improves multimodal reasoning, yielding up to 8.23% accuracy gains over strong baselines and improving BLEU scores up to 8.27% across multiple-choice and open-ended tasks. These findings highlight latent continuous reasoning as a promising direction for advancing LMMs beyond language-bound CoT, offering a scalable framework for human-like reflective multimodal inference.
@@ 1 Introduction
Vision-language models (VLMs) have transformed multimodal tasks, such as visual question answering (VQA), image captioning, and reasoning on benchmarks like ScienceQA [1], MMMU [2], and IQBench [3], by seamlessly integrating visual and textual data. These models leverage visual models and large language models (LLMs) to process heterogeneous inputs, enabling applications from autonomous systems to interactive assistants. However, achieving robust reasoning in VLMs remains a challenge due to limitations in existing techniques, such as attention mechanism within the transformer decoder [4], prompting strategies like CoTs [5, 6, 7], or training methods like reinforcement learning (RL) [8, 9]. CoT, originally developed for LLMs, prompts models to generate intermediate reasoning steps in natural language, while other approaches, such as fine-tuning with visual-text alignment, aim to enhance multimodal reasoning. These methods often rely on discrete token sequences or static vision features, leading to computational inefficiencies and difficulties in dynamically aligning visual and textual modalities for coherent reasoning, particularly in tasks requiring fine-grained multimodal understanding.
Inspired by human cognition, where reasoning involves generating intermediate thoughts and iteratively validating them against input data, such as revisiting images or documents to ensure coherence, we propose the Multimodal Chain of Continuous Thought (MCOUT), a novel framework for efficient reasoning in VLMs. MCOUT operates in a unified latent space, dynamically aligning visual and textual representations to mimic reflective human thinking. We introduce two variants: MCOUT-Base, which uses the language model’s last hidden state as a continuous thought for iterative refinement inspired from COCONUT [10], and MCOUT-Multi, which enhances cross-modal alignment by combining the hidden state with input embeddings via a multimodal latent attention mechanism. By overcoming the limitations of token-based CoT and static vision features, MCOUT reduces computational overhead by directly feeding hidden layers, with/without input embeddings, into the model as continuous thoughts. Tested on diverse benchmarks, MCOUT achieves significant performance gains, positioning it as a pioneering advancement in vision-language reasoning and offering a scalable approach for robust multimodal inference.
@@ 2 Literature Review
The development of reasoning capabilities in VLMs is critical for tasks like VQA and multimodal reasoning. Over the past few years, various techniques have been explored to enhance reasoning in both LLMs and VLMs, including attention mechanism, prompting techniques, training methods, and recent latent reasoning paradigms. CoT prompting, introduced by Wei et al. [5], has significantly improved LLM performance on arithmetic tasks (e.g., GSM8K) and logical reasoning tasks (e.g., AQUA-RAT) by generating explicit intermediate steps. Building on CoT, its variants have emerged to address reasoning limitations. Self-consistency [11] samples multiple CoT outputs and selects the most consistent answer via majority voting, enhancing robustness but increasing computational cost. Tree of Thoughts (ToT) [6] structures reasoning as a tree search, exploring multiple paths for complex problem-solving, though its token-based nature remains computationally intensive. Graph of Thoughts (GoT) [7] extends ToT by modeling reasoning as a graph, enabling dynamic recombination of thoughts for greater efficiency. In the VLM domain, Multimodal CoT [12] generates interleaved text and image reasoning steps, improving performance on ScienceQA but struggling with cross-modal alignment due to reliance on static vision features and verbose token sequences. These CoT methods, while effective for LLMs, face challenges in VLMs, where aligning heterogeneous modalities and minimizing token overhead are critical.
@@
Beyond prompting, training techniques have been pivotal in enhancing reasoning for both LLMs and VLMs. RL methods, such as those explored by Ouyang et al. [8], optimize LLMs using human feedback to improve reasoning and alignment, as seen in models like InstructGPT. Group relative policy Optimization (GRPO) [13] refines model outputs by incorporating reward signals, enhancing performance on complex tasks. Reasoning functions, such as RARL [9], enable models to learn structured reasoning patterns through optimization, improving logical consistency. RL-based fine-tuning [14] has been applied to align visual and textual features, though these methods often rely on static embeddings, limiting dynamic reasoning capabilities. These training techniques complement prompting but still face challenges in efficiently integrating multimodal data for coherent reasoning.
@@
To overcome these limitations, latent reasoning paradigms have shifted reasoning from discrete token sequences to continuous latent spaces. COCONUT [10] leverages the last hidden state of an LLM as a "continuous thought," enabling parallel exploration of reasoning paths via breadth-first search. This approach reduces token overhead and outperforms CoT on tasks requiring backtracking. Other latent reasoning methods for LLMs include Latent Reasoning Skills (LaRS) [15], which uses unsupervised learning to create latent representations of rationales, selecting in-context learning examples based on reasoning skills and achieving fourfold faster processing than CoT. Similarly, Wang et al. [16] proposed a recurrent depth approach that iteratively refines latent representations, scaling test-time computation to enhance performance on complex tasks. These methods demonstrate the efficiency of latent reasoning in LLMs but are primarily designed for text-only contexts, leaving their application to VLMs largely unexplored.
@@
In the VLM domain, latent reasoning is an emerging area with promising developments. Zhang et al. [12] introduced a multimodal CoT framework that uses diffusion processes to learn a text-image aligned latent space, generating dynamic image features that improve reasoning on ScienceQA and multimodal machine translation. Yang et al. [17] developed MMaDA, a diffusion-based VLM that operates in latent spaces for coherent generation and reasoning across text and images, achieving strong performance in tasks like VQA and image captioning. Yang et al. [18] proposed the Mirage framework, which augments VLMs with latent visual tokens during decoding, enhancing reasoning efficiency in complex multimodal tasks. Fan and Zhou [19] introduced stacked latent attention, preserving spatial information in latent spaces to improve reasoning in VQA tasks. Recent efforts, such as multimodal latent language modeling [20], employ next-token diffusion for continuous reasoning, while Corvid [21] and Grounded Chain-of-Thought (GCoT) [22] address visual hallucination and decision-making accuracy. Despite these advances, most approaches rely on discrete token-based reasoning or static vision features, limiting efficient cross-modal alignment.
@@
Inspired by human cognition and previous work [10], where reasoning involves generating intermediate thoughts and iteratively validating them against input data, our MCOUT addresses these gaps by introducing a novel latent reasoning framework for VLMs, specifically for image-based tasks. MCOUT employs two variants: MCOUT-Base, which uses the language model’s last hidden state as a continuous thought for iterative refinement, and MCOUT-Multi, which integrates the hidden state with image embeddings via a multimodal latent attention mechanism, enabling dynamic alignment of visual and textual representations. MCOUT mimics human reflective reasoning by iteratively refining thoughts in a continuous latent space, as demonstrated in our implementation, which supports multimodal inputs and has been tested successfully for vision-language reasoning. MCOUT offers a significant advancement, bridging the efficiency of latent reasoning with the complexity of vision-language reasoning, paving the way for robust and scalable VLMs.
@@ 3.1 Model Architecture
The MCOUT framework is built upon a vision-language model, SilVar [23], comprising a pre-trained visual encoder $\mathcal{V}$ and a language model $\mathcal{L}$, as illustrated in Figure 1. We use CLIP [24] as the visual encoder $\mathcal{V}$, which processes input images $\mathbf{x}_{v}\in\mathbb{R}^{H\times W\times C}$ to produce visual embeddings $\mathbf{e}_{v}\in\mathbb{R}^{S_{v}\times D}$, where $S_{v}$ is the sequence length of visual tokens and $D$ is the embedding dimension. For the language model $\mathcal{L}$, we employ Llama 3.2 1B, which processes tokenized text inputs $\mathbf{x}_{t}$ to generate contextual embeddings $\mathbf{e}_{t}\in\mathbb{R}^{S_{t}\times D}$, where $S_{t}$ is the sequence length of text tokens. In this study, we use CLIP and Llama 3.2 1B for all experiments because we want to focus on latent reasoning for small VLMs, although our pipeline is compatible with other LLMs.
@@ 4.1 Datasets and Training
To evaluate the effectiveness of our MCOUT framework, we conducted experiments using four diverse vision-language datasets: VQAv2 [25], MMMU [2], ScienceQA [1], and MMStar [26]. These datasets assess the model’s reasoning capabilities across multimodal tasks, including VQA, scientific reasoning, and general knowledge understanding, with a focus on image-text integration. The VQAv2 dataset, used for pretraining, contains 443,757 question-answer pairs associated with images from the COCO dataset, emphasizing tasks like object recognition, attribute identification, and spatial reasoning.
@@ 4.2 Results and Benchmarking
Kosmos2 [27] 1.7 32.70
SilVar [23] 7 63.21
LLaVA-7B [28] 7 41.10
InstructBLIP-7B [29] 8 54.10
OpenFlamingo [30] 9 44.80
Qwen-VL [31] 9.6 61.10
MiniGPT-4 [32] 13 47.71
LLaMA2-13B [33] 13 55.78
LLaVA-13B [33] 13 47.74
PandaGPT-13B [34] 13 63.20
@@
Models Parameters (B) accuracy (%) BLEU Our experiments Baseline 1 25.44 25.44 MCOUT-Base ($N_{t}$ = 5) 1 27.53 ($\uparrow$ 8.21%) 27.54 ($\uparrow$ 8.31%) MCOUT-Multi ($N_{t}$ = 5) 1 27.18 ($\uparrow$ 6.79%) 27.19 ($\uparrow$ 6.82%) MCOUT-Base ($N_{t}$ = 10) 1 27.52 ($\uparrow$ 8.18%) 27.54 ($\uparrow$ 8.31%) MCOUT-Multi ($N_{t}$ = 10) 1 27.36 ($\uparrow$ 7.54%) 27.37 ($\uparrow$ 7.58%) Literature reports Kosmos 2 [27] 1.7 23.7 MiniGPT-4-v1-7B [32] 7 23.6 LLaVA-v1.5-7B [28] 7 33.7 MiniGPT-4-v2 [35] 7 25.0 OpenFlamingo v2 [30] 9 28.8 Qwen-VL [31] 9.6 29.6 LLaVA-v1.5-13B [28] 13 37.0 PandaGPT-13B [34] 13 32.9
@@
Models Parameters (B) accuracy (%) BLEU Our experiments Baseline 1 25.13 25.14 MCOUT-Base ($N_{t}$ = 10) 1 26.13 ($\uparrow$ 3.98%) 26.14 ($\uparrow$ 3.98%) MCOUT-Multi ($N_{t}$ = 10) 1 26.07 ($\uparrow$ 3.74%) 26.08 ($\uparrow$ 3.74%) Literature reports Kosmos2 [27] 1.7 24.9 MiniGPT-4-v1-7B [32] 7 16.3 MiniGPT-4-v2 [35] 7 21.3 LLaVA-7B [28] 7 27.1 OpenFlamingo v2 [30] 9 26.9 Qwen-VL-Chat [31] 9.6 34.5 PandaGPT-13B [34] 13 25.6
@@ 4.3 Multimodal Latent Reasoning Analysis
MCOUT-Multi, which integrates the last hidden state with multimodal input embeddings, shows last hidden layer means ranging from -0.02212 to -0.00112 across iterations, with standard deviations around 2.24, closely tracking MCOUT-Base’s patterns and indicating minimal disruption from multimodal fusion. However, the mixed embeddings reveal a critical limitation: their mean remains constant at 0.002418 with a negligible standard deviation of 0.001866, and the mixed standard deviation is uniformly 0.198925 across all iterations, suggesting a static, low-variance contribution from the multimodal input. This persistent uniformity, despite normalization, points to a modality collapse, where the attention mechanism fails to extract diverse visual context, aligning MCOUT-Multi’s performance (58.45% accuracy at $N_{t}=5$) closely with MCOUT-Base (58.60%). This observation resonates with the sinking of visual attention in recent studies [36, 37, 38], which attributes such collapse to activation imbalances favoring a specific type (e.g. text). Our study shows that low-variance embeddings (mixed_embeds std $\approx$ 0.2 vs. last_hidded std $\approx$ 2.2) limit multimodal benefits. The pre-training norm adjustment likely prevented catastrophic fusion failure, but the static mixed embeddings suggest entropy collapse [39], where uniform attention weights diminish multimodal impact.